OpenStudy (anonymous):

An airplane flew for 6 hours against a tail wind of 25 km/hr. The return trip with the same wind took 5 hours. Find the speed of the airplane in still air.

OpenStudy (anonymous):

d= r*t Solve the following for s, the planes speed in no wind.\[6(s-25)=5(s+25) \]\[s=275 \text{ km}/\text{hr} \]

OpenStudy (anonymous):

@shortie1213 Did you close this question?

OpenStudy (anonymous):

no